For a temperature difference ΔT = 20C, one slab of material conducts 10W/m2.


What is the rate of heat flow per m2

of surface area when the slabs are placed side

by side with ΔTtot = 20C?



1
Expert's answer
2021-04-19T07:25:46-0400

For one slab, for a temperature difference ΔT = 20ºC, the rate of heat flow per unit area is 10 W/m2. We know that for one slab "\Delta x" thick, the rate can be expressed as


"q=k\frac{\Delta T_\text{tot}}{n\Delta x}=\frac{20k}{\Delta x}=10\text{ W/K m}^2,\\\space\\
\frac{k}{\Delta x}=0.5."

So, for n slabs


"q_n=k\frac{\Delta T_\text{tot}}{n\Delta x}=\frac{10}{n}\text{ W/m}^2."